# Towards Cross-Lingual LLM Evaluation for European Languages

### v1

@@ Abstract
The rise of Large Language Models (LLMs) has revolutionized natural language processing across numerous languages and tasks. However, evaluating LLM performance in a consistent and meaningful way across multiple European languages remains challenging, especially due to the scarcity of multilingual benchmarks. We introduce a cross-lingual evaluation approach tailored for European languages. We employ translated versions of five widely-used benchmarks to assess the capabilities of 40 LLMs across 21 European languages. Our contributions include examining the effectiveness of translated benchmarks, assessing the impact of different translation services, and offering a multilingual evaluation framework for LLMs that includes newly created datasets: EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K. The benchmarks and results are made publicly available to encourage further research in multilingual LLM evaluation.
Towards Cross-Lingual LLM Evaluation for European Languages
Klaudia Thellmann1 Bernhard Stadler1 Michael Fromm2,3 Jasper Schulze Buschhoff2 Alex Jude2 Fabio Barth4 Johannes Leveling2 Nicolas Flores-Herr2 Joachim Köhler2 René Jäkel1 Mehdi Ali2,3 1TU Dresden, Germany, 2 Fraunhofer IAIS, Germany 3Lamarr Institute, Germany, 4DFKI, Germany thanks: †Equal contribution.
@@ 1 Introduction
The development of Large Language Models (LLMs) has brought transformative advancements to natural language processing, enabling sophisticated tasks such as question-answering, summarization, and machine translation. However, despite these advances, the evaluation of LLMs across languages, especially those beyond English and Chinese, remains a significant challenge. Initiatives like CLEF [^1] https://clef2024.imag.fr/, FIRE [^2] http://fire.irsi.res.in/, and WMT [^3] https://www2.statmt.org/wmt24/ have focused on multilingual evaluations, yet there is a lack of dedicated benchmarks for systematically comparing LLM performance across the languages of Europa.
Creating bespoke benchmarks for each language is a costly and time-consuming endeavor. This not only limits the scalability of evaluations but also results in a fragmented understanding of model performance across different languages. In many cases, evaluations are conducted in isolation, without a basis for meaningful comparison between languages. Moreover, human evaluation is widely recognized as the gold standard for assessing language model performance, as it captures nuances that automated benchmarks may miss. However, human evaluations are resource-intensive and difficult to scale across multiple languages.
To address these challenges, we explore whether automatically translated benchmarks can provide a viable method to estimate human preferences in chat scenarios. While human evaluation would likely be the most accurate way to assess LLM performance, the resource demands of manual dataset creation and evaluation make it impractical for large-scale multilingual comparisons. Instead, this approach aims to reduce these resource demands while ensuring comprehensive evaluation coverage across all languages.
In this paper, we focus on evaluating multilingual LLMs by utilizing translated versions of existing benchmarks. Our goal is to determine the efficacy of translated benchmarks and assess whether they can substitute manually generated ones, allowing for a more scalable and uniform evaluation across multiple languages.
@@
  • We assess whether translated benchmarks reflect human preferences to evaluate their viability in Section 5.
  • We examine how the choice of translation service (i.e., DeeplPro and ChatGPT) affects the performance of current LLMs in accurately responding to various benchmark tasks in Section 6.
  • We release EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K in 20 European languages as evaluation benchmarks for the NLP community [^4] https://huggingface.co/collections/openGPT-X/eu20-benchmarks-67093b13db8ff192dc39312d.
@@ 2 Related Work
Multilingual benchmarks have predominantly been generated using two widely adopted methodologies, besides various intermediate approaches: either by using human annotators of multiple languages to develop a multilingual dataset from scratch Kocmi et al. (2023); Goyal et al. (2022); Conneau et al. (2018) or by using LLMs to translate an existing benchmark into multiple languages Lai et al. (2023); Tiedemann (2012).
Creating a multilingual dataset with human annotators requires a significant amount of time, which impacts both the duration of production and the cost of the dataset Yang et al. (2019). In addition, homogeneous benchmarks are required in each language instead of heterogeneous multilingual benchmarks for a cross-lingual analysis Lewkowycz et al. (2022); Tiedemann (2012).
This method is increasingly inefficient because model scores can become saturated over time Ott et al. (2022), or benchmarks loose their relevance too fast if a desired performance is reached Lewkowycz et al. (2022). An alternative is the machine translation of existing benchmarks. This approach is both more cost-effective and time-efficient. However, the main disadvantage is the translation quality, which is highly dependent on the capability of the language model used Meng et al. (2022). In the past, some data sets have shown quality gaps in translation accuracy, with special consideration given to languages with medium and low resources Team et al. (2022).
Those languages have only a few to very few data sets available for pre-training and evaluation. They are much more difficult to be translated by an LLM, when taking cultural nuances of each language in consideration. Translating these languages, in particular, can lead to imbalanced evaluations across the benchmarks Team et al. (2022); Goyal et al. (2022); Conneau et al. (2018). Therefore, the selection of languages is a central decision to make when creating a multilingual benchmark.
In addition to the choice of languages, it is crucial to consider the range of tasks that need to be covered for a benchmark. For example, creating a multilingual benchmark for a specific task provides valuable insights into model performance for that exact task but says little about a model’s general applicability Conneau et al. (2018). This also includes benchmarks that, for example, cover the question-answering (QA) task for a wide range of domains, as only the QA task is tested here Lewis et al. (2020).
Another approach is to cover multiple tasks to evaluate language models broadly. An example is the OKAPI dataset published by Lai et al. (Lai et al., 2023). For this benchmark Three datasets (AI2 Reasoning Challenge (ARC), HellaSwag, and MMLU) were translated into 26 languages: eight high-resource, 11 medium-resource, and seven low-resource languages Lai et al. (2023).
Building on this approach, we have expanded the linguistic reach through additional translations with special consideration on European languages and increased the variety of tasks for a wider capability test of LLMs. Furthermore, we show that our data pre-processing has improved quality in certain language areas (See Section 3).
@@ 3.1 Translation Process
We chose DeepL as translation service for its balance between translation accuracy and scalability. We translated five well-known datasets, ARC Clark et al. (2018), HellaSwag  Zellers et al. (2019), TruthfulQA Lin et al. (2022), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2020) from English into 20 European languages. These datasets encompass a mix of multiple-choice and open-ended generation tasks, each of which presented unique translation challenges. The translations maintained the original structure of each task to ensure the benchmarks were as consistent as possible across languages. Using DeepL’s XML-tag handling, we avoided issues with prompt formatting and ensured that key contextual elements, such as ’Question’ and ’Response’ prompts, were properly localized (cf. Table 5).
To enhance the visibility and accessibility of our benchmarks, we launched a multilingual leaderboard [^5] https://huggingface.co/spaces/openGPT-X/european-llm-leaderboard on Hugging Face, inspired by the popular Open LLM leaderboard [^6] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard.
@@ 3.2 Evaluation Process
We conducted the model evaluations on the HPC Cluster Alpha Centauri (TUD/ZIH) [^7] https://compendium.hpc.tu-dresden.de, the JUWELS Booster (JSC) [^8] https://apps.fz-juelich.de/jsc/hps/juwels/index.html, and, for larger models, on the JURECA Cluster (WEST AI/JSC) [^9] https://westai.de/services/hardware/, using systems powered by NVIDIA A100 and H100 Tensor Core GPUs. For model evaluation, we used the LM Evaluation Harness Framework from Eleuther.AI [^10] https://github.com/OpenGPTX/lm-evaluation-harness in versions v0.4.1 and v0.4.3, evaluating in a multi-GPU setting using hf/accelerate with both data and model parallelism. For larger models, we deployed a local vllm-based inference server (v0.5.3) [^11] https://github.com/vllm-project/vllm/releases/tag/v0.5.5 using the LM Evaluation Harness local completions implementation (commit 42dc24/main) to perform pipeline-parallel evaluations across multiple GPUs on a single node.
4 European LLM Evaluation Results
This section provides an overview of the performance of large language models (LLMs) across various sizes, tasks and languages. Section 4.1 focuses on the evaluation across all 21 European languages and indicates which models provide a competitive performance in a multilingual setting. Section 4.2 shows the connection between the proportion of languages in the CommonCrawl dataset and the performance of the current best open weights models. Lastly, Section 4.3 highlights how the performance of the models depend on the language family. Given that many models provide little to no information about the composition of their training corpora, we focus our detailed analysis on the LLMs that demonstrate the most reliable performance across 21 languages.
4.1 Global European LLM Evaluation
This section provides an overview of the performance of several LLMs on various tasks and languages, grouped by their number of parameters in small-sized ($\leq$9B parameters), medium-sized ($>$9B & $<$27B parameters) and large ($>$27B parameters) models. The models are evaluated on our newly introduced tasks EU21-ARC, EU21-GSM8K, EU21-HellaSwag, EU21-MMLU, and EU21-TruthfulQA. An extensive overview of all models are provided in the Appendix as Tables (cf. LABEL:tab:acc_std_21lang_1, LABEL:tab:acc_std_21lang_2) and task specific heatmaps (cf. Figure 5, 6 7, 8, 9).
@@ 4.1 Global European LLM Evaluation
Among the small-sized models, the focus is on Gemma-2-9b-Instruct  Team et al. (2024b), Meta-Llama-3.1-8B-Instruct  Dubey et al. (2024), Mistral-NeMo-Minitron-8B-Base Sreenivas et al. (2024), and Qwen2-7B-Instruct Yang et al. (2024), further models with a lower average performance like Gemma-7b Team et al. (2024a) are in the appendix LABEL:tab:acc_std_21lang_1 and LABEL:tab:acc_std_21lang_2
Gemma-2-9b-Instruct consistently performs well (cf. Figure 1), by having a narrow interquartile ranges, signaling low variability and therefore stable performance in all languages. On the multilingual grade school math word benchmark EU20-GSM8K the median accuracy is worse than Meta-Llama-3.1-8B-Instruct, indicating lower math capabilities. In EU20-MMLU the model shows a robust behaviour, but the median is slightly lower than Qwen2-7B-Instruct. Both latter benchmarks indicate that the capacity of small models might not allow for both a reliable performance on all languages, aswell as the capability to handle specialized knowledge reliable.
Meta-Llama-3.1-8B-Instruct also delivers solid results, occasionally surpassing Gemma-2-9b-Instruct, such as in the EU21-GSM8K task where it achieves the highest median performance. While capable in math related tasks, the model shows shortcomings across all 21 languages compared to Gemma-2-9b.
Mistral-NeMo-Minitron-8B-Base performs competitively in certain tasks, particularly in more widely spoken languages, indicated by competitive maximum scores. However, its overall performance is weaker compared to other two models, with significant drops in less-resourced languages and complex tasks. In the EU21-GSM8K and EU21-MMLU tasks the median performance is the worst compared to the other models.
@@
This analysis (cf. Figure 2) focuses on medium-sized LLMs such as Mistral-Nemo-Instruct-12.2B [^12] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024), Phi-3-medium-14B-4k-Instruct Abdin et al. (2024) and Vicuna-13b-v1.5 Peng et al. (2023).
Mixtral-8x7B-Instruct-v0.1 shows the highest median performance on four out of the five tasks. Only on EU21-GSM8K its performance is lacking behind Mistral-Nemo-Instruct-12.2B, indicating worse performance on math related tasks. However, its performance tends to be more variable than Mistral-Nemo-Instruct-12.2B_2407, with some languages like Latvian displaying significantly lower scores, reflecting higher inconsistency across the board.
Mistral-Nemo-Instruct-12.2B_2407 delivers solid results across all tasks and languages, with a particularly strong median performance in the EU21-GSM8K task. As evidenced by its narrow interquartile ranges the model signals low variability across all 21 languages compared to the other models.
Phi-3-medium-14B-4k-Instruct presents competitive results mostly in English by having the highest GSM8K, MMLU values but lags vastly behind the previous two models in median peformance across all languages and tasks, indicating only a small amount of multilingual training corpora. Lastly, Vicuna-13-b-v1.5 based on the older LLama-2-13B is not competitive on any task, due to its missing multilingual capabilities.
Compared to the small-sized category, Mixtral-8x7B-Instruct-v0.1 provides a substantial gain in EU21-Hellswag by achieving an median performance of 0.64, an absolute increase of 4% compared to Gemma-2-9b-Instruct. On EU21-GSM8K Mistral-Nemo-Instruct-12.2B achieves 57% accuracy a small increase of 1% compared to Meta-Llama-3.1-8B-Instruct. On EU21-MMLU Qwen2-7B achieved a accuracy of 59.4%, slightly behind medium sized models such as Mistral-Nemo-Base-12.2B with 60.2% and Mixtral-8x7B-Instruct-v0.1 with an performance of 60.8%.
@@
This analysis (cf. Figure 3) focuses on large-sized LLMs such as Gemma-2-27b-Instruct Team et al. (2024b), Vicuna-33b-v1.3 Peng et al. (2023), Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) and c4ai-command-r-35B-v01131313https://huggingface.co/CohereForAI/c4ai-command-r-v01. Gemma-2-27b-Instruct and Meta-Llama-3.1-70B-Instruct reliable outperforms the other models across most tasks, with median scores often exceeding an accuracy of 0.7. Their performance is not only high but also stable across different languages, as evidenced by its narrow interquartile ranges, signaling low variability. Probably due to the higher capacity (compared to Gemma-2-9b), Gemma-2-27b-Instruct can now also score higher on specialized knowledge tasks such as MMLU and GSM8K.
Vicuna-33b-v1.3 and c4ai-command-r-35B-v01 perform worse compared to the previous two models. This indicates a monolingual training corpora paired with a monolingual tokenizer.

### v2

# Towards Multilingual LLM Evaluation for European Languages

@@ Abstract
The rise of Large Language Models (LLMs) has revolutionized natural language processing across numerous languages and tasks. However, evaluating LLM performance in a consistent and meaningful way across multiple European languages remains challenging, especially due to the scarcity of language-parallel multilingual benchmarks. We introduce a multilingual evaluation approach tailored for European languages. We employ translated versions of five widely-used benchmarks to assess the capabilities of 40 LLMs across 21 European languages. Our contributions include examining the effectiveness of translated benchmarks, assessing the impact of different translation services, and offering a multilingual evaluation framework for LLMs that includes newly created datasets: EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K. The benchmarks and results are made publicly available to encourage further research in multilingual LLM evaluation.
Towards Multilingual LLM Evaluation for European Languages
Klaudia Thellmann1 Michael Fromm2,3 Bernhard Stadler1 Jasper Schulze Buschhoff2 Alex Jude2 Fabio Barth4 Johannes Leveling2 Nicolas Flores-Herr2 Joachim Köhler2 René Jäkel1 Mehdi Ali2,3 1TU Dresden, Germany, 2 Fraunhofer IAIS, Germany 3Lamarr Institute, Germany, 4 DFKI, Germany thanks: †Equal contribution.
@@ 1 Introduction
The development of Large Language Models (LLMs) has led to transformative advancements to natural language processing, enabling complex tasks such as question-answering, summarization, and machine translation. However, despite these advances, the evaluation of LLMs across languages—especially those beyond English and Chinese —remains a significant challenge. Initiatives like CLEF [^1] https://clef2024.imag.fr/, FIRE [^2] http://fire.irsi.res.in/, and WMT [^3] https://www2.statmt.org/wmt24/ have focused on multilingual evaluations, yet there is a lack of dedicated benchmarks that allow for consistent and comparable evaluation of LLM performance across different European languages.
Creating custom benchmarks for each language is a costly and time-consuming, which limits the scalability of evaluations and can result in a fragmented understanding of model performance across different languages. Without comprehensive multilingual evaluations, comparisons between languages are often constrained, leading to an incomplete understanding of how models perform in languages beyond high-resource ones like English, German, or French.
To address these challenges, we investigate whether automatically translated benchmarks can reliably assess model performance across languages and whether this performance correlates with human evaluations, similar to benchmarks conducted in English.
In this paper, we focus on evaluating multilingual LLMs using translated versions of existing benchmarks. Our goal is to determine the effectiveness of these translated benchmarks and assess whether they can substitute manually generated ones, allowing for a more scalable and uniform evaluation across multiple languages.
@@
  • We assess whether translated benchmarks correlate with human evaluations to asses their viability in Section 5.
  • We compared translation services in Section 6, in particular DeeplPro and ChatGPT, to identify correlations in LLM performance and understand how translation quality impacts the models’ benchmark results.
  • We release EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K in 20 European languages as evaluation benchmarks for the NLP community. [^4] https://hf.co/collections/openGPT-X/eu20-benchmarks-67093b13db8ff192dc39312d.
@@ 2 Related Work
Multilingual benchmarks have predominantly been generated using two widely adopted methodologies: either through human annotation to create datasets from scratch Kocmi et al. (2023); Goyal et al. (2022); Conneau et al. (2018), or by translating existing benchmarks using large language models (LLMs) Lai et al. (2023); Tiedemann (2012).
While datasets developed by human annotators offer high-quality translations and task-specific accuracy, they require significant time and financial investment Yang et al. (2019). Furthermore, as models improve, these benchmarks can quickly become outdated, limiting their usefulness Ott et al. (2022); Srivastava et al. (2023).
In contrast, the machine translation of existing benchmarks is more efficient and cost-effective, but its quality is often inconsistent, especially for medium- and low-resource languages, where translation inaccuracies can impact model evaluation Meng et al. (2022); NLLB Team et al. (2022).
These translation issues can lead to imbalanced evaluations across languages, particularly in languages with limited pre-training and evaluation resources Goyal et al. (2022); Conneau et al. (2018). Thus, selecting the right languages and ensuring translation quality is critical for creating a robust multilingual benchmark.
In addition to the choice of languages, it is crucial to consider the range of tasks that need to be covered for a benchmark. Creating a multilingual benchmark for a specific task provides valuable insights into model performance for that exact task but says little about a model’s general language capability Conneau et al. (2018). This also includes benchmarks that, for example, cover the question-answering (QA) task for a wide range of domains, as only the QA task is tested Lewis et al. (2020).
Another important factor is the range of tasks covered in a benchmark. Task-specific benchmarks provide insights into a model’s performance for specific tasks but do not fully capture a model’s overall language capabilities Conneau et al. (2018); Lewis et al. (2020). To address this, benchmarks that span multiple tasks, such as the OKAPI dataset Lai et al. (2023), offer a broader evaluation of model capabilities. OKAPI, for example, translates four diverse tasks (ARC, HellaSwag, MMLU and TruthfulQA) into 31 languages, including high-, medium-, and low-resource languages.
Building on this approach, we have expanded the linguistic reach through additional translations with special consideration on European languages and extend the variety of tasks. By including GSM8K Cobbe et al. (2021) we offer evaluations on multilingual grade school math word problems. Furthermore, we show that our data pre-processing has improved quality in certain language areas (See Section 3).
@@ 3.1 Translation Process
We chose DeepL as translation service due to its balance between translation accuracy and scalability. We translated five well-known datasets, ARC Clark et al. (2018), HellaSwag  Zellers et al. (2019), TruthfulQA Lin et al. (2022), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2020) from English into 20 European languages. These datasets encompass a mix of multiple-choice and open-ended generation tasks, each presenting unique translation challenges. The translations preserved the original structure of each task to ensure consistency across languages. By leveraging DeepL’s XML tag handling, we avoided issues with prompt formatting and ensured that key contextual elements, such as ’Question’ and ’Response’ prompts, were appropriately localized (cf. Table 7). To enhance the visibility and accessibility of our benchmarks, we launched a multilingual leaderboard [^5] https://hf.co/spaces/openGPT-X/european-llm-leaderboard on Hugging Face, inspired by the popular Open LLM leaderboard [^6] https://hf.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard.
@@ 3.2 Evaluation Process
We conducted the model evaluations on the HPC Cluster Alpha Centauri (TUD/ZIH) [^7] https://compendium.hpc.tu-dresden.de, the JUWELS Booster (JSC) [^8] https://apps.fz-juelich.de/jsc/hps/juwels/index.html, and, for larger models, on the JURECA Cluster [^9] https://westai.de/services/hardware/, using systems powered by NVIDIA A100 and H100 Tensor Core GPUs. For model evaluation, we used adapted versions of the LM Evaluation Harness Framework from Eleuther.AI in versions v0.4.1 and v0.4.3, evaluating in a multi-GPU setting using hf/accelerate with both data and model parallelism. For larger models, we deployed a local vLLM-based inference server (v0.5.3) [^10] https://github.com/vllm-project/vllm/releases/tag/v0.5.5 using the LM Evaluation Harness local-completions implementation111111commit 42dc24 LM Evaluation Harness to perform pipeline-parallel evaluations across multiple GPUs on a single node. We expended about 45,000 GPU hours conducting the model evaluations.
4 Evaluation Results
This section provides an overview of the performance of LLMs across various sizes, tasks and languages. Section 4.1 focuses on the evaluation of all 21 European languages considered in this paper and investigates which models provide a competitive performance in a multilingual setting. Section 4.2 explores the relationship between the proportion of languages in the CommonCrawl dataset and the performance of the best current open-weights model. Lastly, Section 4.3 examines how model performance varies across different language families. Since many models disclose little to no information about the composition of their training corpora, we focus our detailed analysis on the LLMs that exhibit the most consistent performance across the 21 languages.
4.1 European LLM Evaluation
This section provides an overview of the performance of several LLMs on various tasks and languages, grouped by their number of parameters in small-sized ($\leq$9B parameters), medium-sized ($>$9B & $<$27B parameters) and large ($>$27B parameters) models. The models are evaluated on our newly introduced tasks EU21-ARC, EU21-GSM8K, EU21-HellaSwag, EU21-MMLU, and EU21-TruthfulQA. A comprehensive overview of all models are provided in the Appendix, including Tables (cf. LABEL:tab:acc_std_21lang_1, LABEL:tab:acc_std_21lang_2) and task specific heatmaps (cf. Figure 5,  6,  7,  8,  9).
@@ 4.1 European LLM Evaluation
Among the small-sized models, the focus is on Gemma-2-9b-Instruct  Team et al. (2024b), Meta-Llama-3.1-8B-Instruct  Dubey et al. (2024), Mistral-NeMo-Minitron-8B-Base Sreenivas et al. (2024), and Qwen2-7B-Instruct Yang et al. (2024). Models with a lower average performance, like Gemma-7b Team et al. (2024a), are listed in Appendix Tables LABEL:tab:acc_std_21lang_1, LABEL:tab:acc_std_21lang_2.
Gemma-2-9b-Instruct performs consistent across languages (cf. Figure 1), as indicated by its narrow interquartile range, but lags behind Meta-Llama-3.1-8B-Instruct on the EU21-GSM8K benchmark, signaling lower math abilities. In EU21-MMLU, the model shows a robust behaviour, but the median is slightly lower than Qwen2-7B-Instruct. The results on the latter two benchmarks indicate that the capacity of small models might not allow for a reliable performance on all languages and a good performance on specialized knowledge.
Meta-Llama-3.1-8B-Instruct excels in math-related tasks, achieving the highest median performance in EU21-GSM8K, but its performance lags behind Gemma-2-9b-Instruct across all 21 languages.
Mistral-NeMo-Minitron-8B-Base shows competitive performance in widely spoken languages but struggles in less-resourced languages and complex tasks. Its performance on EU21-GSM8K and EU21-MMLU is notably lower than the other models.
@@
This analysis (cf. Figure 2) focuses on medium-sized LLMs such as Mistral-Nemo-Instruct-12.2B [^12] https://hf.co/mistralai/Mistral-Nemo-Instruct-2407, Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024), Phi-3-medium-14B-4k-Instruct Abdin et al. (2024) and Vicuna-13b-v1.5 Peng et al. (2023).
Mixtral-8x7B-Instruct-v0.1 shows the highest median performance on four out of the five tasks. Only on EU21-GSM8K its performance is lacking behind Mistral-Nemo-Instruct-12.2B. However, its performance tends to be more variable across languages than Mistral-Nemo-Instruct-12.2B_240, indicated by the low minima in 4 out of 5 tasks.
Mistral-Nemo-Instruct-12.2B_2407 delivers solid results across all tasks with a particularly strong median performance in the EU21-GSM8K task. As evidenced by its narrow interquartile ranges, the model signals low variability across all 21 languages compared to the other models.
Phi-3-medium-14B-4k-Instruct presents competitive results in English by having the highest GSM8K and MMLU values, but lags vastly behind the previous two models in median peformance across all languages and tasks, indicating a focus on a few selected languages in the training corpora. Lastly, Vicuna-13-b-v1.5 based on the older LLama-2-13B is not competitive on any task, due to its limited two trillion English-centric training tokens.
Compared to the small-sized category, Mixtral-8x7B-Instruct-v0.1 provides a substantial gain in EU21-HellaSwag by achieving an median performance of 0.64, an absolute increase of 4% compared to Gemma-2-9b-Instruct. On EU21-GSM8K Mistral-Nemo-Instruct-12.2B achieves 57% accuracy a small increase of 1% compared to Meta-Llama-3.1-8B-Instruct. On EU21-MMLU Qwen2-7B achieved a accuracy of 59.4%, slightly behind medium sized models such as Mistral-Nemo-Base-12.2B with 60.2% and Mixtral-8x7B-Instruct-v0.1 with an performance of 60.8%.
@@
This analysis (cf. Figure 3) focuses on large-sized LLMs such as Gemma-2-27b-Instruct Team et al. (2024b), Vicuna-33b-v1.3 Peng et al. (2023), Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) and c4ai-command-r-35B-v01131313https://hf.co/CohereForAI/c4ai-command-r-v01.
Gemma-2-27b-Instruct and Meta-Llama-3.1-70B-Instruct reliable outperforms the other models across most tasks, with median scores often exceeding an accuracy of 0.7. Their performance is not only high but also stable across different languages, as evidenced by its narrow interquartile ranges, signaling low variability. We assume that due to the higher capacity (compared to Gemma-2-9b), Gemma-2-27b-Instruct now also scores higher on specialized knowledge tasks such as EU21-MMLU and EU21-GSM8K.
Comparable, to the smaller Vicuna model also the 33 billion parameter version is not competitive on any task, due to its limited two trillion English-centric training tokens. c4ai-command-r-35B-v01 performs better in English compared to the Vicuna model.